quadratic equation roots are -1 and 2, leading coefficient 4
Answer by TimothyLamb(4379)

(x + 1)(x - 2) = 0
xx - 2x + x - 2 = 0
xx - x - 2 = 0
4(xx - x - 2 = 0)
4xx - 4x - 8 = 0
---
Free algebra tutoring live chat:
https://sooeet.com/chat.php?gn=algebra
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations with quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php
---